Please help me graph the line that satisfies each condition:
1. slope=-1, contains P(-2, 1)
2. contains Q(-1, 3) and is perpendicular to line AB with A(-2, 0) and B(4, 3)
3. contains M(1, -1) and is parallel to line FG with F(3, 5) and G(-3, -1)
4. slope=-3/4, contains K(3, -2)

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

1. slope=-1, contains P(-2, 1)
----
m = slope
Use y-y1 = m*(x-x1) where (x1,y1) is the point.
===================
2. contains Q(-1,3) and is perpendicular to line AB with A(-2,0) and B(4,3)
Find the slope, m1, of AB
The slope of lines perpendicular, m2, is -1/m1
The it's the same as #1
========================
3. contains M(1,-1) and is parallel to line FG with F(3,5) and G(-3,-1)
---
Find the slope, m, of FG.
The slopes of parallel lines are the same.
etc.
======================
4. slope=-3/4, contains K(3,-2)
y-y1 = m*(x-x1)
*********************************************************
That gives you the equation of each line.
they're straight lines, so you need 2 points for a graph.


Answer by ikleyn(53938)

.

In this site, there is a group of lessons devoted to similar problems. These lessons are

    - Find the slope of a straight line in a coordinate plane passing through two given points
    - Equation for a straight line having a given slope and passing through a given point
    - Solving problems related to the slope of a straight line
    - Equation for a straight line in a coordinate plane passing through two given points
    - Equation for a straight line parallel to a given line and passing through a given point
    - Equation for a straight line perpendicular to a given line and passing through a given point (*)

    - OVERVIEW of lessons related to the slope of a straight line

So, if you want to learn the subject to the level to solve such problems on your own, you can do it reading my lessons.

Consider them as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.